OpenStudy (anonymous):

Both angle XAB and angle XYZ are right angles, XB = 6, BY = 2 and AB = 4. The ratio of the area of triangle XAB/ area of triangle XYZ is (a) 3/5 (b) 5/16 (c) 9/25 (d) 4/5 (e) 3/25

OpenStudy (anonymous):



OpenStudy (anonymous):

Need the picture to do this... :\

OpenStudy (callisto):

Use ratio of similar figures... triangle XAB is similar to triangle XYZ (AAA) By Pyth thm, \(XA=\sqrt{XB^2 - AB^2} = \sqrt{6^2 - 4^2}=\sqrt{20}\) ratio of areas = (ratio of sides)^2 So, \[\frac{\text{area of triangle XAB}}{\text{area of triangle XYZ}} \ = \ (\frac{XA}{XY})^2\]\[\frac{\text{area of triangle XAB}}{\text{area of triangle XYZ}} \ = \ (\frac{\sqrt{20}}{6+2})^2\]Time for you to work out the answer!